### 09 Dec 2023 v1

# History Matters: Temporal Knowledge Editing in Large Language Model

Abstract
The imperative task of revising or updating the knowledge stored within large language models arises from two distinct sources: intrinsic errors inherent in the model which should be corrected and outdated knowledge due to external shifts in the real world which should be updated. Prevailing efforts in model editing conflate these two distinct categories of edits arising from distinct reasons and directly modify the original knowledge in models into new knowledge. However, we argue that preserving the model’s original knowledge remains pertinent. Specifically, if a model’s knowledge becomes outdated due to evolving worldly dynamics, it should retain recollection of the historical knowledge while integrating the newfound knowledge. In this work, we introduce the task of Temporal Knowledge Editing (TKE) and establish a benchmark AToKe (Assessment of TempOral Knowledge Editing) to evaluate current model editing methods. We find that while existing model editing methods are effective at making models remember new knowledge, the edited model catastrophically forgets historical knowledge. To address this gap, we propose a simple and general framework termed Multi-Editing with Time Objective (METO) for enhancing existing editing models, which edits both historical and new knowledge concurrently and optimizes the model’s prediction for the time of each fact. Our assessments demonstrate that while AToKe is still difficult, METO maintains the effectiveness of learning new knowledge and meanwhile substantially improves the performance of edited models on utilizing historical knowledge.
1 Introduction
Large-scale language models (LLMs) have made impressive progress in the last few years (Brown et al. 2020; Ouyang et al. 2022; Touvron et al. 2023; OpenAI 2023). However, the knowledge in a language model always needs to be updated because the internal knowledge of the model can be wrong and the external world knowledge is constantly changing (Sinitsin et al. 2020; Hartvigsen et al. 2022; Ji et al. 2023). It would be costly to retrain the model each time, so there is a lot of work proposing knowledge editing (KE) methods that allow new correct knowledge to be injected directly into specific model parameters. Previous work includes constrained fine-tuning (Zhu et al. 2020), hypernetwork knowledge editing (Cao, Aziz, and Titov 2021; Hase et al. 2021; Mitchell et al. 2022a), external memory-based editing (Mitchell et al. 2022b; Zhong et al. 2023; Zheng et al. 2023) and locate-then-edit model editing (Dai et al. 2022; Meng et al. 2022b). All of these methodologies focus on making the model memorize new knowledge.
Figure 1: An example of an edited GPT-J model losing historical knowledge by using the existing editing method MEMIT. The editing operation overwrites Joseph Biden being the President of the U.S. against Donald Trump, but does not preserve historical knowledge, which causes Trump’s relationship with the U.S. President to be lost.
However, a clear differentiation is essential during the process of knowledge editing, specifically discerning between two distinct scenarios: (1) knowledge correction, involving rectification of inaccurate knowledge arising from the model’s training data which needs to be corrected, and (2) knowledge updating, necessitated by shifts in the external world or evolving cognitive paradigms which needs to be updated. Existing work on KE lacks this crucial distinction, and is committed to making the LLMs memorize the new knowledge, while simply ignoring the original knowledge in LLMs. However, within the context of knowledge updating, we believe that retaining historical knowledge within the model holds great value. For example, as shown in Figure 1, although we are all “updated” with the knowledge that President of the United States is Joseph Biden, we still sometimes want to know who the former President was.
Figure 2: Presentation of Temporal Knowledge Editing Task. Suppose that the training corpus for the model is collected in 2016, so the internal time of the model is 2016. After one edit the model can obtain the knowledge that Donald Trump is the president of the United States in 2017 and Obama is the former president. Then keep editing in this way, the model’s internal time keeps moving forward while not losing history and ensuring temporal ordering.
The current evaluation of KE methods focuses on new knowledge being memorized and irrelevant knowledge being left unaltered, while not considering the appropriate keeping of relevant historical knowledge. Therefore, we propose the task of Temporal Knowledge Editing (TKE) and construct a benchmark AToKe for evaluating KE methods in historical knowledge preservation, by collecting series of world knowledge with timestamps and viewing them as the form of a series of knowledge updates. We explore single editing, multiple editing and extending editing (e.g., letting the model know that the president is still Biden in 2023 in Figure 2) of facts on the time series. After each editing, we ask questions about historical knowledge and current knowledge.
We evaluate state-of-the-art knowledge-editing methods on AToKe and find that they are effective in making the model memorize new knowledge but cause confusion in time, e.g., in our example, the knowledge about the former president disappeared in edited LLMs.
In order to address this problem to some extent and to enhance the effectiveness of current model editing methods, we propose a simple and general framework, METO, which edits both historical and new knowledge and optimizes the model’s prediction for the time of each fact.
Our contributions are listed below:
  1. 1.
    We categorize knowledge editing scenarios, point out the issue that current methods corrupt historical knowledge, and validate it through experiments and analysis.
  2. 2.
    We are the first to propose the Temporal Knowledge Editing task, and publish a benchmark AToKe for evaluating the task.
  3. 3.
    We propose a new editing framework METO that improves the performance of previous editing methods on the benchmark.
Our code and dataset will be released to the community.
2 Temporal Knowledge Editing
This section introduces our arguments about knowledge editing as well as definitions and evaluation for the temporal knowledge editing task.
2.1 Background: Knowledge in Language Models
There has been a lot of work considering LLMs as knowledge bases and accessing, applying and manipulating the knowledge in them (Petroni et al. 2019a; AlKhamissi et al. 2022). Previous work always represents knowledge as a triple $(s,r,o)$, consisting of a subject ($s$), a relation ($r$), and an object ($o$) (e.g. (U.S., president, Joseph Biden)) and explores how to extract knowledge from LLMs (Bordes et al. 2013; Lin et al. 2015). In this paper, we follow the work that uses discrete prompt to extract knowledge (Petroni et al. 2019b; Davison, Feldman, and Rush 2019). Specifically, we construct a natural language template $q_{r}(\cdot)$ for each relation $r$, which is then combined with a subject $s$ in a knowledge triple to generate a question or a cloze-style statement $q_{r}(s)$, which is also consistent with the previous work of the knowledge editing (Meng et al. 2022a; Zhong et al. 2023). It is a natural way to extract or test the knowledge in a language model.
2.2 Requirements for Knowledge Editing
What are the requirements that a good knowledge editing method needs to satisfy? We believe that the edited model should be changed in some ways and hold constant in others.
Aspects of Change
The knowledge we edit is naturally expected to change in the model, which is generally referred to as Edit Success in prior work in evaluation. In addition, the paraphrase expression about the knowledge should also be changed, which is referred to as Paraphrase Success in evaluation. Similarly, the knowledge associated with the edited knowledge should also be changed, which has been explored by Zhong et al. (2023) and referred to as Multi-hop Accuracy.
Aspects of Constancy
Except for what is relevant to the edited knowledge, it seems like no other knowledge or ability of the model should be changed. Previous work (Meng et al. 2022b) has defined Neighborhood Success to evaluate whether other knowledge of the original entity is affected. They also define Reference Score to check if generations from the edited model are semantically consistent with the new object. And Generation Entropy is designed to test ability of the edited model to generate fluent sentences.
However, no work has yet explored the knowledge to be edited. We argue that if the knowledge within the model is outdated (changed to historical knowledge) due to the flow of time, then the historical knowledge should also be preserved within the model. It is quite reasonable because if not then the timeline in the model would be messed up and our former president would disappear in the aforementioned example. Therefore, we propose the task of Temporal Knowledge Editing (TKE) [^1] Temporal Knowledge Editing focuses on knowledge that is outdated due to the flow of time and therefore needs to be edited, rather than original errors within the model..
2.3 Task Definition
As mentioned in Section 2.1, we use triples to represent knowledge. In addition, we attach temporal scope to each relational fact $\{(s,r,o,t_{s},t_{u})\}$ since the relational fact often shows temporal dynamics, where $t_{s}$ and $tu$ denote the fact $(s,r,o)$ valid from $ts$ to $tu$. As time passes, the same subject and relation will correspond to different objects in sequence, such as (U.S., president, Donald Trump, 2017, 2021), (U.S., president, Joseph Biden, 2021, N/A) and so on. Consistent with previous work, we use $(s,r,o,t_{s},t_{u})\rightarrow(s,r,o^{\star},t_{s}^{\star},t_{u}^{\star})$ to represent an knowledge editing operation, denoted as $e$, meaning that the object of subject $s$ under relation $r$ changes from $o$ to $o^{\star}$. And the reason for the edit is that $(s,r,o)$ is valid from $t_{s}$ until $t_{u}$, and after that $(s,r,o^{\star})$ is valid from $t_{s}^{\star}$ to $t_{u}^{\star}$ [^2] Note that the $t_{u}$ is vacant if the knowledge is still valid currently.. Specifically, $o$ and $o^{\star}$ can be the same object, at which point the editing operation $(s,r,o,t_{s},t_{u})\rightarrow(s,r,o,t_{s},t_{u}^{\star})$ will make the model know that the fact $(s,r,o)$ lasts until $t_{u}^{\star}$, where $t_{u}^{\star}$ is later than $t_{u}$. For convenience of expression, we refer to $(s,r,o,t_{s},t_{u})$ as historical knowledge and $(s,r,o^{\star},t_{s}^{\star},t_{u}^{\star})$ as current knowledge.
Given a collection of knowledge editing operations =$\mathcal{E}=\{e_{1},e_{2},\dots\}$ and a language model $\mathcal{M}$, knowledge editor $F$ aims to get an edited model $\mathcal{M}_{e}=F(\mathcal{M},\mathcal{E})$ that satisfy the requirements mentioned in Section 2.2. Specifically, in temporal knowledge editing, for the edited model $\mathcal{M}$, we expect that it should not only memorize the current knowledge $(s,r,o^{\star})$, but the knowledge in it should also be organized on the timeline. In other words, the model can answer correctly when asked questions about the “previous” or “last one” knowledge $(s,r,o)$, or when asked directly about the relevant knowledge $(s,r,?)$ at a certain moment in time between $t_{s}$ and $t_{u}$. For example, after updating the model’s knowledge of U.S. president to Biden, the test questions include “Who was the former president of the United States?” or “Who was the president of the United States in 2005?” .
2.4 Evaluation of Temporal Knowledge Editing
Since time is constantly moving forward, a piece of knowledge may be updated multiple times in different temporal scopes. Therefore, we define three evaluation subtasks: 1) Temporal Knowledge Single Editing (TSE), which requires that, after a single update to a piece of knowledge, the model sustains temporal chronology. 2) Temporal Knowledge Multiple Editing (TME), demanding ordered temporal succession after multiple unidirectional knowledge updates. It is worth pointing out the difference between TME and mass-editing of prior work (Meng et al. 2022b; Mitchell et al. 2022a): mass-editing refers to the simultaneous editing of multiple facts, which are usually unrelated, while TME refers to the continuous editing of the same $s$ and $r$ at different time scopes. 3) Temporal Knowledge Extending Editing (TEE), which requires that duration of target knowledge perceived by the model is successfully prolonged. More details on the benchmark are presented in Section 3.
3 AToKe: Assessment of Temporal Knowledge Editing
We introduce the AToKe (Assessment of temporal knowledge editing) benchmark, which comprises three distinct datasets as outlined in Table 1. These datasets are specifically designed to evaluate the effectiveness of knowledge editing methods in handling temporal information. The first dataset, referred to as AToKe-SE, consists of temporal knowledge Single Edits. The second dataset, AToKe-ME, includes temporal knowledge Multiple Edits associated with the same subject and relation. The third dataset, AToKe-EE, encompasses knowledge Edits that Extend the temporal scope of the original knowledge. In the following sections, we first present the collection of the basic temporal knowledge and further describe how the three datasets were extracted from it. We then detail the data statistics and evaluation settings of the datasets, and finally present the evaluation metrics.
BTK (United States, head of government,
(Obama, 2009, 2017), (Trump, 2017, 2021),
(Biden, 2021, 2022))
SE (United States, head of government,
(Obama, 2009, 2017) → (Trump, 2017, 2021))
[…]
(Obama, 2009, 2017) → (Trump, 2017, 2021)
→ (Biden, 2021, 2022))
[…]
(Biden, 2021, 2022) → (Biden, 2021, 2023))
$\mathcal{Q}$ a) Who is the current President of the United States? (ALL)
b) Who was the President of the United States
in the previous term? (SE&ME)
c) Who holds the position of President in the United States
from 2017 to 2021? (SE&ME)
d) Who was the President of the United States
from 2009 to 2017? (SE&ME)
e) From 2022 to 2023, who serves as the president
of the United States? (EE)
$\mathcal{A}$ a) Donald Trump
b) Barack Obama
c) Donald Trump
d) Barack Obama
e) Joseph Biden
Table 1: An instance illustrating the construction of the dataset, where the collected base temporal knowledge (BTK) serves as the foundation. The dataset encompasses three types of edits, namely single edit (SE), multiple edits (ME), and extending edit (EE), along with corresponding questions ($Q$) and their respective answers ($A$) following the knowledge editing. In this case, the model has been edited from Obama to Trump.
3.1 Basic Temporal Knowledge Collection
Sampling time series facts
Our dataset is based on YAGO3.0.3333https://yago-knowledge.org/downloads/yago-3 (Mahdisoltani, Biega, and Suchanek 2015), a knowledge base comprising fact triples associated with millions of entities extracted from Wikipedia, enriched with WordNet, GeoNames, and other data sources. Following the previous work (Dasgupta, Ray, and Talukdar 2018), we begin by sampling temporal facts from the YAGO data. We extract all fact triples with their time scope and obtain the temporal facts set $\{(s,r,o,t_{s},t_{u})\}$. Subsequently, after matching by $s$ and $r$ and sorting by $t_{s}$ with all these facts, we obtain the chain of temporal facts $\mathcal{C}=\{(s,r,o_{1},t_{s_{1}},t_{u_{1}}),...,(s,r,o_{N},t_{s_{N}},t_{u_{N}})\}$. Finally, to ensure non-overlapping and non-contradictory time sequences within the chain, we employ a heuristic algorithm. See the Appendix A.1 for more technical details.
Locating Time of Model Facts
For a chain of temporal facts $\{(s,r,o_{N},t_{s_{N}},t_{u_{N}})\}$, its $o$, $t_{s}$ and $t_{u}$ are constantly changing. Therefore, to test the performance of knowledge editing, we need to determine where the knowledge is located in the model. First, we use the GPT-J model [^4] Without loss of generality, we use GPT-J as the representative LLM used in the experiments. to verify and filter out $(s,r,*)$ that the model does not have. Next, we locate the model’s knowledge position in the temporal fact chain. By asking questions about the facts in the temporal chain, we find the most recent fact in each chain that the model has successfully recalled as the model’s initial knowledge. See Appendix A.1 for more details of locating.
Sampling Future Fake facts
So far, we have collected a temporal chain of facts from the YAGO dataset where the latest knowledge is collected in 2022. To ensure that there is always new knowledge used to update for all models in our dataset and to make our dataset effective to be used as the benchmark for TKE in the long run, we design future fake facts to augment the original temporal chain. Specifically, for each chain, we sample one counterfactual object in YAGO to serve as counterfactual object $o_{N+1}$ that is in the same class as real objects. Then $o_{N+1}$ will be randomly assigned a reasonable time scope and appended to the end of the chain. Finally, we obtain a temporal chain of facts that incorporates constructed fake facts.
Generating Temporal natural language questions
As mentioned in Section 2.1, given a chain of temporal facts that we construct, we construct template $q_{r}(\cdot)$ for each relation $r$. The constructed questions are categorized into two categories based on time: 1) Explicit time question, which refers to asking the question that have an explicit time frame. 2) Relative time question, on the other hand, does not contain an explicit time frame and uses words like “present”, “previous” or “last one” instead. We first utilize ChatGPT (gpt-3.5-turbo) to automatically generate questions for all relations, and then manually filter them. The details of question generation such as prompts used, question examples and statistics of questions are shown in Appendix A.1.
3.2 Construction of Three Datasets
Up until now, we have obtained the basic temporal knowledge, including temporal fact chains, questions and answers at different time points (respectively BTK, Q and A in Table 1). And after locating and expanding, the first fact of every temporal fact chain represents the current knowledge of the model. Therefore, all of our edits begin with the first fact. Moving forward, we will construct three temporal knowledge editing datasets.
AToKe-SE
For Single Edit (SE), we take the first two facts from each chain in BTK to construct a single edit operation, and ask questions about both historical and current facts.
AToKe-ME
Multiple edits (ME) are superimposed on single edits. We similarly traverse all temporal chains, construct successive editing operations for all facts in each chain, and ask questions about facts with each time scope.
AToKe-EE
Extending Edit (EE) is an edit operation that extends the time scope of the current fact of the model. We traverse all chains of temporal facts and select the first fact in each chain as the extending object. By manually extending this object, we keep the target entity $o$ unchanged but extend the time scope. This edit operation simulates the situation where certain facts are not changed in the future.
Method AToKe-SE AToKe-ME AToKe-EE
Current Historical Current Historical Edited Current
CES CES-P CRS HES HRS CES CES-P CRS HES HRS $*{}^{*}$ CES CES-P CRS
CFT 5.73 5.69 5.34 0.06 0.02 1.11 1.18 1.22 0.03 0.00 0.01 3.41 2.91 3.15
MEND 80.47 40.56 32.46 1.73 0.68 71.83 27.96 23.67 0.40 2.10 0.25 91.94 62.48 51.63
ROME 99.99 97.01 81.64 2.41 1.56 98.85 91.54 77.08 0.44 1.17 0.26 99.93 98.70 85.84
MEMIT 99.66 92.23 75.31 2.22 1.21 98.42 91.06 66.48 0.48 0.86 0.27 99.92 95.82 72.76
Table 2: Results of existing models on the benchmark AToKe-SE, AToKe-ME and AToKe-EE. “Edited” means the score is computed when all the multiple edits are completed. The best results are highlighted in BOLD.
3.3 Dataset Summary
Dataset format
The three datasets have a similar data format, but they differ in the number of edits and the questions used to evaluation. These datasets all contain edit sets (SE&ME&EE), questions ($\mathcal{Q}$), standardized answers to questions ($\mathcal{A}$), and answer alias sets. In these datasets, both single edit and multiple edits contain questions about historical and current knowledge, while extending edits contain only questions about the current knowledge. Single edit and extending edit have only one editing operation and corresponding questions for one $(s,r)$, while multiple edits contain multiple operations and questions. Format examples for the three datasets can be found in the Appendix A.2
Data statistics
In the base temporal data collection phase, we first sample 169,996 temporal fact chains. After filtering by GPT-J, there are 13 different relations left and the number of our temporal chains is reduced to 8,820, which is the final number of the three datasets. For the length distribution of temporal fact chains, fact chains of length 2-5 account for more than 90%, with very few chains exceeding 10. In addition, the relationship ”playsFor” has the largest number of fact chains. The time is measured in years. And for the time scope, before adding the future fake facts, the latest knowledge is obtained in 2022, and after that it is extended to 2028. More detailed statistics are in the Appendix A.2.
3.4 Evaluation Metrics
As mentioned before, we categorize questions into relative time questions and explicit time questions, and both of them are about historical and current knowledge. Therefore, we have the following four indicators to assess the performance of TKE: 1) Historical Relative time Question Score (HRS), which is the accuracy of relative time questions about historical knowledge answered by the edited model. 2) Historical Explicit time Question Score (HES), which is the accuracy of explicit time questions asked about historical knowledge. 3) Current Relative time Question Score (CRS), which is the accuracy of relative time questions about current knowledge. 4) Current Explicit time Question Score (CES), which measures accuracy of explicit time questions about current knowledge. In addition, the prompt we used as the optimization target when editing is the same as the questions in CES. Therefore, to make results more valid, we use paraphrases of the CES questions for further evaluation, denoted as 5) Current Explicit time Paraphrase Question Score (CES-P).
In particular, in AToKe-ME, we treat each edit as a single edit and compute the score as in AToKe-SE, and finally report the average of all edits. In addition, we calculate HES for all the historical facts at the end of multiple consecutive edits of the model to measure the overall editing performance (denoted as $*{}^{*}$). In AToKe-EE, since there is no historical knowledge, we do not need to measure HRS and HES. CThe results of detailed formal definitions and another metric can be found in the Appendix A.3.
4 Current Status on TKE
In order to better assess the performance of existing methods for temporal knowledge editing, we conduct experiments with several popular editing methods and analyze the results.
4.1 Experimental Setup
Knowledge Editing Methods
We evaluate the following state-of-the-art editing methods on AToKe. 1) Constrained Fine-tuning (CFT) (Zhu et al. 2020) performs gradient descent on the target knowledge to minimize the loss and set a norm constraint on model weight changes. 2) MEND (Mitchell et al. 2022a) learns a hypernetwork to produce weight updates using a low-rank decomposition of the gradient obtained by standard fine-tuning. 3) ROME (Meng et al. 2022a) firstly localizes the factual knowledge in the model by causal tracing, and then treats the MLP modules as key-value stores to insert new knowledge by making a rank-one change. 4) MEMIT (Meng et al. 2022b), a successor to ROME, can insert lots of memories at once by modifying the MLP weights of a range of critical layers.
Figure 3: Demonstration of the METO editing framework. First the model will be queried based on the current knowledge to get the knowledge under the model time ($C_{m}$). Then both historical and current knowledge are used as target knowledge for target knowledge optimization and time objective optimization with any model editing methods.
All of the above model editing methods make changes to the parameters of the model, which is the focus of our exploration. In addition, there are some methods based on external storage that do not edit the model. The performance of that class of methods depends on the knowledge classifiers or retrievers, which is analyzed and explored in the Appendix B.
Given a knowledge editing operation that is expected to be learned $e=(s,r,o,t_{s},t_{u})\rightarrow(s,r,o^{\star},t_{s}^{\star},t_{u}^{\star})$, we convert it to a cloze statement by natural language template $q_{r}(\cdot)$ which is used as the input to the above knowledge editing approaches. Note that all the previous methods use only $q_{r}((s,r,o))$ as input; to ensure that the information is sufficient and adapted to our task, we have expanded these methods to include time for each input cloze statement.
Language Model to be Edited
Following setup of previous work (Meng et al. 2022a, b; Zhong et al. 2023), we use GPT-J (6B) (Wang and Komatsuzaki 2021) as the base LLM to be edited with above methods.
4.2 Results and Analysis
The performance of existing knowledge editing methods on our benchmarks is shown in Table 2. For models, CFT is weaker and hence performs poorly. ROME performs slightly better than MEMIT, probably because we edit one piece of knowledge at a time and hence are better suited to ROME. After comparison and analysis we can get the following main conclusions:
Remembering the new and forgetting the old
Except for CFT, edited models can remember new knowledge very well and do some generalization, but the performance on historical knowledge is disastrous. It is consistent with our expectations, as all existing model editing methods optimize the probability of current knowledge and ignore historical knowledge. An example is shown in Figure 1, where Donald Trump is forgotten by the edited models.
Relative time questions are more difficult than explicit time questions
We can observe that it is more difficult for the model to answer questions about a piece of knowledge without explicitly providing it with a specific time, using a relative time expression such as “last one”. It may be because using relative time expression requires the model to reason about the order in which facts occur, whereas using explicit time expression simply requires the model to remember when facts occur.
The more edits, the worse the performance
The average results on AToKe-ME are sightly worse than those on AToKe-SE, which may be because editing multiple times on the same fact causes a little confusion to the edited model. As we can see that HES score is only about 0.2% after multiple edits, and all the knowledge injected before the last edit is invalidated.
Extending time scope is easier than inserting new object
We can see that methods on AToKe-EE perform the best of three datasets, which is reasonable because the extending edit task does not change the current knowledge, but merely extends its time scope. The correct answer is not changed when the question is asked about the fact.
5 METO: Multi-Editing with Time Objective
As we can see from the previous experiments in Section 4.2, the existing methods perform excellently in memorizing new knowledge, but catastrophic forgetting of historical knowledge occurs after editing. To alleviate this serious problem, we propose a simple editing framework METO (Multi-Editing with Time Objective) that can be applied to enhance existing editing methods easily.
Method AToKe-SE AToKe-ME
Current Historical Current Historical Edited
CES CES-P CRS HES HRS CES CES-P CRS HES HRS $*{}^{*}$
${}^{+}$ 2.8↓2.93 2.62↓3.07 2.26↓3.08 3.38↑3.32 2.43↑2.41 1.27↑0.16 1.2↑0.02 1.04↓0.18 1.64↑1.61 1.23↑1.23 0.73↑0.71
${}^{+}$ 83.26↑2.79 33.45↓7.11 25.41↓7.05 30.14↑28.41 30.17↑29.49 70.52↓1.31 28.41↑0.45 21.84↓1.83 28.65↑28.25 30.83↑28.73 21.83↑21.58
${}^{+}$ 99.95↓0.04 93.78↓3.23 78.88↓2.76 20.25↑17.84 16.29↑14.73 99.93↑1.08 90.97↓0.57 82.4↑5.32 23.22↑22.78 18.18↑17.01 15.92↑15.66
${}^{+}$ 86.4↓13.26 85.32↓6.91 74.07↓1.24 30.31↑28.09 24.32↑23.11 92.73↓5.69 85.75↓5.31 73.58↑7.10 36.2↑35.72 26.04↑25.18 21.93↑21.66
Table 3: Results of enhanced models with METO (marked with “${}^{+}$”) on the benchmark AToKe. Changes of the results are labeled on the side. The best results are highlighted in BOLD.
5.1 Methodology
As shown in Figure 3, We firstly query the language model with the cloze statement of current knowledge to extract the knowledge under model time, and then edit the model using both of them with timestamps, and also use the model to make knowledge time predictions so that the model can reinforce both historical and new knowledge, and improve the awareness of time of knowledge.
Model-time Knowledge Extraction
For a chain of temporal facts $\mathcal{C}=\{(s,r,o_{1},t_{s_{1}},t_{u_{1}}),...,(s,r,o_{N},t_{s_{N}},t_{u_{N}})\}$, the object is constantly changing and the knowledge of the model about $(s,r,\cdot)$ is determined by the collection time of its training corpus. Therefore, in order to better preserve the historical knowledge inside the model, we need to extract the model-time knowledge about $(s,r,\cdot)$. With the method in Section 3.1, we obtain the fact of what is happening under the model time which is noted as $C_{m}=\{(s,r,o_{i},t_{s_{i}},t_{u_{i}})\}$. By comparing $\mathcal{C}$ and $C_{m}$, we can get the knowledge that needs to be newly captured by the model $C_{m+}=\{(s,r,o_{i+1},t_{s_{i+1}},t_{u_{i+1}}),...,(s,r,o_{N},t_{s_{N}},t_{u_{N}})\}$.
Multi-editing on Both Historical and Current Knowledge
Combining $C_{m}$ and $C_{m+}$, we obtain the set of all the target knowledge $C_{t}$ which is used as the target of editing. Using the target knowledge $C_{t}$ obtained above as input, we can edit the language model with any model editing method, such as MEND, ROME, MEMIT, etc. Note that in the specific implementation, to suit our task, we have also added timestamps in the cloze statement of original methods.
It is worth explaining that we only use $C_{m}$ which is occurring at the model time and do not use the full previous knowledge, $C_{m-}$, as the historical knowledge. It is because we believe that the model already knows that $C_{m-}$ is the history that has happened in the past which does not need to be changed. For $C_{m}$, on the other hand, the model needs to be made aware of the duration of this knowledge, which has ended and changed from the present tense to the past tense.
Time Objective Optimization
In order to further enhance the model’s awareness of the time of knowledge, along with the multi-editing described above, we perform an additional task of optimizing the time objective of knowledge. We also use the same editing method of existing model to ensure the generalization of our framework, except that the editing target is changed from the object to the corresponding time. As an example, given an input of “Donald Trump is the President of the United States from”, the model is then edited to optimize the probability of “2017 to 2021” with ROME or other methods.
5.2 Results on AToKe-SE and AToKe-ME
Since existing methods perform relatively well on AToKe-EE and it does not involve the questioning of historical facts, we test with METO enhancement for editing on AToKe-SE and AToKe-ME (shown in Table 3).
We can find that maintaining little change in performance on current knowledge, our framework greatly improves existing editing methods’ performance on historical knowledge. Among them, ROME stills performs best in memorizing current knowledge. Surprisingly, MEND and MEMIT perform better on historical knowledge, which may be due to the fact that our framework prefers that editing methods can edit more than one piece of knowledge at a time, which in turn can preserve historical knowledge.
The editing experiments with our framework also further validate our two previous conclusions: 1) relative time questions are more difficult than explicit time questions. 2) the more edits, the worse the performance. It is worth noting that both phenomena have been mitigated to some extent.
It is promising that, as shown in Table 3, the $*{}^{*}$ score on AToKe-ME improves from the original result of less than 0.3 to more than 15 (except for CFT). Such a significant improvement in this most difficult metric also shows that our framework is beneficial and the edited model remembers some of the previously injected knowledge.
Although there has been a substantial improvement in memorizing historical knowledge, it is still far from a satisfactory level, reflecting the difficulty of our proposed task of temporal knowledge editing, which still requires a concerted effort by the community.
6 Related Work
The expanding parameter count of language models leads to higher retraining costs. And since the knowledge inside a model becomes progressively outdated, knowledge editing (KE), a convenient way to edit the knowledge, has received increasing attention and some methods have been proposed. Zhu et al. (2020) propose the constrained finetuning approach on modified facts to solve the problem. Sotoudeh and Thakur (2019) utilize symbolic representations and generalized RELU networks. (Hahnloser et al. 2000) to correct model with small patches. Cao, Aziz, and Titov (2021) introduce the method which corrects knowledge and improves predictions using a hyper-network for targeted modifications. Dai et al. (2022) explore knowledge neurons in models and attempt to leverage them to edit specific factual knowledge. ROME (Meng et al. 2022a) and MEMIT (Meng et al. 2022b) take the approach of locating and then editing. Mitchell et al. (2022a) a collection of small auxiliary editing networks that use a single desired input-output pair to make fast, local edits. In addition, MEMIT-CSK (Gupta et al. 2023) is designed to edit commonsense knowledge in GPT.
Evaluation metrics on KE generally focus on whether the editing is successful and whether other unrelated knowledge is affected. Onoe et al. (2023) also evaluate abilities of updated LLM by making inferences based on injected facts. Furthermore, Zhong et al. (2023) build multi-hop questions to assess edited models on related facts. And Cohen et al. (2023) evaluate the ripple effects in the edited models. However, no work has yet noted that pre-editing historical knowledge should also be preserved, which we argue is very important.
7 Conclusion
In this paper, we systematically classify scenarios involving knowledge editing, identify the shortcomings leading to historical knowledge distortion in existing model editing methods. To facilitate a comprehensive evaluation of KE techniques, we introduce the task of temporal knowledge editing (TKE) and present a new benchmark named AToKe. We conduct experiments on AToKe and demonstrate that existing methods lead to a catastrophic forgetting of historical knowledge. To bridging existing gaps, we present the METO editing framework, enhancing the efficacy of preceding approaches. However, TKE still remains challenging and calls for more efforts in the community.
Appendix A Appendix A
A.1 Details of Dataset Construction
Sampling temporal fact chains from yago
We use the YAGO version 3.0.3 (Dasgupta, Ray, and Talukdar 2018) open source dataset that contains files for all aspects of the knowledge base content. We mainly use the ”yagoMetaFacts.tsv” and ”yagoFacts.tsv” files to build the temporal knowledge editing dataset. ”yagoMetaFacts.tsv” file has the format of [Fact_Time_ID, Fact_ID, Start or End Timestamp, Start or End Time]. Based on this file, we sample to get all the facts and their start or end dates. ”yagoFacts.tsv” file is in the format of [Fact ID, Fact Subject, Fact Relationship, Fact Object]. We locate specific fact triples in the second file based on the fact ID in the first file. With these two files, we can get all the triples containing fact dates: (fact subject, fact relation, fact object, fact start or end time), i.e., ${(s,r,o,t_{s}\,or\,t_{t})}$. Among them, there are 4325485 triples of factual occurrences and 1145969 triples of factual endings. The number is different because there are some events that have occurred but have not yet ended. After matching, we end up with 1145969 fact triples containing time intervals, forming the set of temporal facts $\{(s,r,o,t_{s},t_{u})\}$. Finally, we matched all the facts and formed a chain of triples with the same $s$ and $r$ in the order of the start time of the facts, obtaining a chain of 169996 temporal facts.
Upon observation, we found that there are some overlapping time intervals in the current temporal chain. To simplify the current dataset [^5] Of course, we could choose to keep all time intervals, but this would increase the complexity of the task. We plan to consider including the fact of overlapping time intervals in subsequent releases., we consider removing overlaps and conflicts in these time intervals. We adopt the following three principles to remove conflicting events: 1) Cover the timeline as broadly as possible to cover more time periods, 2) Retain as many event counts as possible, 3) Prioritize events with shorter time intervals.
Locating Facts Time in GPT-J
We use GPT-J to locate the model fact time in the temporal fact chain. If all facts in a temporal chain cannot be recalled, we filter the temporal chain directly. If more than one fact in the temporal chain is successfully recalled, we use the latest successfully recalled fact as the model fact time and remove all facts before that fact. If only one fact recall is successful, we locate that time as the time of the model fact and delete all facts prior to that fact. For each relation, we define question templates to ask about each fact, and the questions are of ”completion” form. Table 4 below shows an example of the relation playfor.
playfor
(In-context-learning examples)
1. Luke Shaw plays for Manchester United.
2. Luka Modrić plays for Real Madrid.
3. Mohamed Salah plays for Liverpool.
4. Wardell Stephen Curry plays for Golden State Warriors.
5. Patrick Lavon Mahomes plays for Kansas City Chiefs.
Completion: Lionel Messi plays for
Answer: Barça.
Table 4: Examples of prompt we use to recall facts with the relation playfor. We use in-context-learning with 5 demonstration examples to ensure that the model outputs answers in the desired format.
Generating Questions using ChatGPT
System:
You are a very useful assistant and your task is to generate relation-related questions as follows:
User:
Given a relation (<graduatedFrom>) provided by the user, you need to generate a question related to that relation. Here, the relation (rel) refers to the connection between two entities in a knowledge graph. Specifically, if we have A <bornIn> B, and the user provides the input <bornIn>, you should consider A as the known information and ask a question about B. Therefore, the generated question should be: Where was [TMP] born?
Examples:
User input: <playFor>
Your output: In which team is [TMP] currently a player?
[normal-[[[4 in-context demonstrations abbreviated]normal-]]]
Hint: Use [TMP] as a placeholder for the entity.
System:
"What university did [TMP] graduate from?"
"From which university did [TMP] graduate?"
"From which institution did [TMP] graduate?"
Table 5: An example of using ChatGPT to generate questions. We manually write 5 demonstrations as the prompt when querying ChatGPT.
We use ChatGPT to construct templates of questions for each relationship and to check them manually. We list the templates used to ask questions to ChatGPT in Table 5.
A.2 Dataset Summary
Data Format
SE (United States, head of government,
(Obama, 2009, 2017) → (Trump, 2017, 2021))
$\mathcal{Q}$ CES: Who holds the position of President
in the United States from 2017 to 2021?
[…]
the United States from 2017 to 2021?
[…]
in the United States?
[…]
the United States from 2009 to 2017?
[…]
$\mathcal{A}$ Donald Trump
Donald Trump
Donald Trump
Barack Obama
Barack Obama
ME (United States, head of government,
(Obama, 2009, 2017) → (Trump, 2017, 2021)
→ (Biden, 2021, 2022))
$\mathcal{Q}$ CES: 1. Who holds the position of President
in the United States from 2017 to 2021?
 2. Who holds the position of President
in the United States from 2021 to 2023?
[…]
the United States from 2017 to 2021?
 2. Who currently serves as the president of
the United States from 2021 to 2023?
[…]
President in the United States?
 2. ***
[…]
the United States from 2009 to 2017?
 2. Who was the President of
the United States from 2021 to 2023?
[…]
 2. ***
$\mathcal{A}$ 1. Donald Trump 2. Joseph Biden
1. Donald Trump 2. Joseph Biden
1. Donald Trump 2. Joseph Biden
1. Barack Obama 2. Donald Trump
1. Barack Obama 2. Donald Trump
EE (United States, head of government,
(Biden, 2021, 2022) → (Biden, 2021, 2023))
$\mathcal{Q}$ CES: Who holds the position of President
in the United States from 2017 to 2023?
[…]
the United States from 2017 to 2023?
[…]
$\mathcal{A}$ Barack Obama
Barack Obama
Barack Obama
Table 6: The case of three datasets AToKe-SE, AToKe-ME and AToKe-EE. In ME, 1 and 2 denote the first and second edits in multiple edits. *** denotes the same as the previous question.
Cases for the three datasets can be found in Table 6.
Data Statistics
Figure 4: Length distribution of temporal fact chains
Relation Temporal Chain Num Avg. Length
playsFor 7358 3.32
isMarriedTo 779 2.05
isAffiliatedTo 438 2.16
graduatedFrom 124 2.06
created 37 2.08
hasWonPrize 38 2.03
worksAt 17 2.35
owns 25 2.00
isLeaderOf 3 2.00
Table 7: Number of temporal chains for each relation and their average lengths
Figure 5: Schematic of the time evolution of AToKe-SE and AToKe-ME. The first color change from the bottom of the bar graph indicates AToKe-SE. x denotes the chain of temporal facts, here we sampled 50 chains.
Figure 6: Schematic of the time evolution of AToKe-EE. x denotes the chain of temporal facts, here we sampled 50 chains.
As shown in Figure 4, we can observe that more than half of the chains in the temporal fact chain have a length of 2. This suggests that sampling future fake facts to expand the data is essential. In addition, only a very small fraction of the chains have a length of more than 10. Overall, the number of chains decreases as the length of the chain increases. Table 7 shows the length distribution in terms of relationships. Most of the chains belong to the first relation playfor and this relation contains chains of multiple lengths. All other relations have chains of less than 1000. Figures 5 and 6 show the time variation of the three datasets.
A.3 Evaluation Metrics
We evaluate the editing methods based on five metrics: HRS, HES, CRS, CES and CES-P. These five metrics are used to compute the accuracy of the results on five different questions.
Suppose we have edited a language model $LM^{*}(\cdot)$ and the probability of answering $o$ to a question is denoted as: $\mathbb{P}\left[LM^{*}(s,r,q)=o\right]$. Given an editing operation $e=(s,r,o\to o^{*})$, the accuracy rate of our metrics is
$\mathds{1}\{\mathbb{P}\left[LM^{*}(s,r,q)=o^{*}\right]>\mathbb{P}\left[LM^{*}(s,r,q)=o^{{V}}\right]\}$
where $o^{V}$ is all the words that can be decoded from the vocab.
In addition, Meng et al. (2022a) uses different metrics from ours such as Efficacy Score(ES) and Paraphrase Scores(PS) for measuring the efficacy of significant changes. The expression for Efficacy Score is:
$\mathds{1}\{\mathbb{P}\left[LM^{*}(s,r,q)=o^{*}\right]>\mathbb{P}\left[LM^{*}(s,r,q)=o\right]\}$
where $o$ is the entity before the model is edited.
It is important to note here that our metrics require that the edited results must answer the question correctly, i.e., the model decodes a result that agrees with the golden reference answer. The latter metrics, on the other hand, only require that the probability of the post-edited result is greater than the probability of the pre-edited result.
CES CRS HES HRS
Single Edit 97.56 84.58 98.51 33.01
Multiple Edits 95.85 70.64 95.98 22.65
Extending Edit 97.75 91.75 N/A N/A
Table 8: Result of the Memory-Based Method
Appendix B Appendix B
Exploring Memory-Based Method for Temporal Knowledge Editing
In the body paragraphs, this paper mainly discusses temporal knowledge editing methods that modify parameters to change the model’s behavior. Another possible type of knowledge editing method is based on retrieval-augmented generation, which means retrieving relevant facts from an edited knowledge base and asking the model to use both the original query and the retrieved facts for reasoning, hoping that the model can produce the desired output. However, these methods rely heavily on the model’s reasoning and in-context learning abilities, and when the model itself is not capable enough, these methods will be much less effective than methods based on modifying parameters. Moreover, the performance of retrieval-augmented generation methods is sensitive to the quality of the prompt and the retriever. Recently, advanced large language models, such as ChatGPT, have demonstrated remarkable in-context learning abilities, which enable retrieval-augmented generation methods to perform better on these models. This has inspired the exploration of knowledge editing methods based on retrieval-augmented generation. Inspired by model editing methods based on memory (Zhong et al. 2023; Zheng et al. 2023), we provide a retrieval-augmented generation method for temporal knowledge editing. The method utilizes the gpt-3.5-turbo model, which has superior reasoning capabilities, as the base model and is evaluated on the AToKe benchmark that we introduce. Despite its simplicity, the method proves to be effective for editing the knowledge of large language models. We consider it a valuable addition to this study, as well as a reference for future research on temporal knowledge editing.
This method does not try to alter the internal knowledge of the model but rather uses a retrieval-reasoning process to modify the answer generated by the model, thus achieving knowledge editing. Specifically, when the language model answers a question, it first obtains the corresponding timestamp as the question time based on the current question and then generates a preliminary answer based on the time and question. Next, the process uses a knowledge retriever (we use Contriver (Izacard et al. 2021) as the knowledge retriever in this method) to retrieve the top 5 facts with timestamps that are most relevant to the question. These facts are concatenated with the preliminary answer generated by the model and fed back to the language model for answering the question. Based on the context examples and Chain of Thought (Wei et al. 2023) guidance that we provide, the language model will perform reasoning, that is, select the facts that are needed, and compare them with the answer generated by the model. If the answer generated by the model is found to be contradictory to the updated facts, the model will then use the updated facts as the final answer and return it. In addition, since different subtasks in AToKe have different question formats, we also use different context examples and CoT methods for different subtasks to improve the performance of the model.
An possible example of the reasoning process is shown in table 9.
Assuming the current time is 2012, answer the following questions.
Question: Which league is Aleksandar Vučić affiliated to?
Generated answer: Aleksandar Vučić is affiliated to Serbian Radical Party.
Retrieved fact:
  From 2008 to 2012, Aleksandar Vučić is affiliated to Ankara
  From 1993 to 2008, Aleksandar Vučić is affiliated to Serbian Radical Party
  From 2008 to 2011, Igor Mirović is affiliated to Ankara
  From 2007 to 2008, Slobodan Vuković is a player of FK Palić
  From 2000 to 2004, Slobodan Vuković’s team is Red Star Belgrade
The current time is 2012, so the retrieved fact that should be selected is “From 2008 to 2012, Aleksandar Vučić is affiliated to Ankara”
Retrieved fact contradicts to generated answer, so the intermediate answer is: Ankara
Final answer: Ankara
Table 9: Reasoning Process of the Memory-Based Method
Table 8 shows the results of the provided temporal knowledge editing method based on memory retrieval. It can be observed that benefitting from gpt-3.5-turbo’s powerful context learning and reasoning capabilities, the method performs well on CES and HES. However, it is worth noting that, even though we provide the current time as an extra input when answering questions, the language model still has low accuracy on CRS; furthermore, when querying for HRS, the model fails to select the most suitable fact from the retrieved facts in most cases, resulting in lower accuracy, i.e., 33.01 for Single Edit HRS and 22.65 for Multiple Edits HRS. The results suggest that AToKe is still challenging even for the powerful gpt-3.5-turbo with memory-based temporal knowledge editing. We hope that this result can inspire future research and promote more applications of temporal knowledge editing.
Thanks for your careful and valuable comments. We will explain your concerns point by point.
B.1 Response to Reviewer #2
Answer for Q1: memory-based editing baselines
As shown in Table 10, we evaluate MeLLo on AToKe (using gpt-3.5). The results of MeLLo suggest that despite the strong reasoning and contextual learning abilities of gpt-3.5, it still underperforms in recalling historical knowledge. We will add results of MeLLo and SERAC in the final version.
Task CES CRS HES HRS
Single Edit 84.69 83.67 21.43 19.39
Multiple Edits 71.43 73.54 14.81 14.29
Extending Edit 90.75 91.25 N/A N/A
Table 10: Results of MeLLo on AToKe
Answer for Q2: results on larger models
Our results fall into two categories, one is Section 4.2, which points out shortcomings of baselines (denoted as res1), and the other is Section 5.3, which points out the improvements brought by METO (denoted as res2). We believe that res1 will not change because baselines are working to optimize the learning of current knowledge only. Therefore, models cannot recall historical knowledge well even they become larger. As for res2, we believe that METO can better capture current and historical knowledge as models get larger and the base editing method gets stronger. Thus, the dimensions of res2 will further become better.
Answer for Q3: impact of each component
As shown in Table 2 here and Table 3 in the paper, we breakdown the impact of ”Multi-Editing” (ME) and ”Time Objective Optimization” (TOO) on the results. We observe that both ME and TOO effectively enhance performance.
AToKe-ME Current Historical Edited
CES CES-P CRS HES HRS HRS
${}^{+}$ 1.27 1.2 1.04 1.64 1.23 0.73
w/o TOO ↓0.03 ↓0.01 ↑0.02 ↓0.36 ↓0.20 ↓0.21
${}^{+}$ 70.52 28.41 21.84 28.65 30.83 21.83
w/o TOO ↑0.27 ↓0.07 ↑0.30 ↓7.97 ↓7.21 ↓3.87
${}^{+}$ 99.93 90.97 82.4 23.22 18.18 15.92
w/o TOO ↓0.32 ↑0.15 ↓0.57 ↓3.95 ↓3.54 ↓2.30
${}^{+}$ 92.73 85.75 73.58 36.2 26.04 21.93
w/o TOO ↑0.86 ↑1.59 ↓0.77 ↓8.62 ↓5.94 ↓6.08
Table 11: The performance of the model after removing the key and value of our design successively.
B.2 Response to Reviewer #3
Answer for Weaknesses 1: limitations and challenges of METO
METO, as the framework we propose to augment previous baseline methods, presents simplicity and universality. At the same time it also has some limitations and challenges. 1) Computational Resource: As you mentioned in Questions 1, multiple edits are performed in order to memorize both historical and current knowledge (HK and CK). It does increase the computational resource consumption of METO. Detailed analysis will be provided in Answer for Q1. 2) Knowledge Distortion: As shown in Table 3 in the paper, the METO-enhanced baselines show a decrease in performance on CK (1-5%), although there is a very significant improvement in performance on HK (around 20%). This suggests that METO introduces some knowledge distortion. However, we believe that this cost is worthwhile and a trade-off. 3) Impact of Base Editing Methods: Since METO is an augmentation framework, the knowledge distortion problem depends more on base editing methods. As shown in Figure 3, the gains from MEMIT and MEND, which are more applicable to multiple editing, are more significant (around 30%). Therefore, the application of METO needs to consider the capabilities of base editing methods.
Answer for Weaknesses 2: accuracy-based metrics
In practice, all previous knowledge editing works rely on accuracy-based metrics. However, we totally agree with you that it might benefit from more diverse evaluation metrics to capture different aspects. We also hope that the evaluation metrics for knowledge editing can be refined in future work.
Answer for Weaknessses 3: future research directions
Please see Answer for Weakness 1 and Answer for Q3.
Answer for Q1: computational resources
METO framework does introduce additional computational requirements. For convenience, we denote methods that only can edit one knowledge at a time as NM, such ROME, and the opposite as MM, such as MEMIT. For NM, with METO for enhancement, the space cost remains the same and the time cost becomes 2n/(n-1) times the original. For MM, with METO for enhancement, the time cost is unchanged and the space cost becomes 2n/(n-1) times of the original. n is the number of times of multiple edits. Due to space constraints, we will add proof in the final version. Thank you very much.
Answer for Q2: time objective optimization
(refer to Figure 3 in the paper) Our Time Objective Optimization: optimize all knowledge for temporal information. For example: for (U.S., President, Biden, ? , ?) , we will optimize 2021-2023. In other words, we edit the time of knowledge into the model using base editing methods. It enhances the model’s ability to memorize the temporal information of knowledge, thus better retaining historical information. Referring to the supplementary experiment in the reply to Review2 Q3, we demonstrate the effect of TOO through ablation study.
Answer for Q3: future research directions
First, in terms of model interpretability, we believe that the temporal awareness of the model deserves further measurement and exploration, such as the relationship between temporal knowledge being stored inside the model. Secondly, in terms of knowledge editing methods, it is also worth investigating how to make the model aware of temporal changes more naturally so that historical knowledge can be properly placed. Finally, in terms of knowledge editing (KE) metrics, we propose to include temporal KE as one metric. This means that KE methods are required to satisfy localization and accuracy while being able to handle changes with related knowledge. It contributes to the development and improvement of KE